The perimeter of a rectangular outdoor patio is 76 ft. The length is 4 ft greater than the width. What are the dimensions of the patio?
Answer #1

Let length = x

Hence, width =x-4

Perimeter = 2(l+w) =76

or, 2x-4 = 38

or, x= 21

Hence, length = 21 ft, width = 17 ft